Don’t use excel

3. X and Y are 2 random variables described by the following regression model: The following random sample of 3 is drawn: i (x,Y) 2 (5,12) 3 (12,20) a. Calculate b, what is the meaning of the number you calculated for ? c. Write down the sample regression line. d. Is it appropriate to use the sample regression line to predict the value of Y when X-10? If your answer is no, why is it inappropriate? If your answer is yes, what is the regressions prediction for the value of Y when X= 10? e. Find e f. Draw a sketch of the sample regression line and put in 3 dots representing the 3 sample observations. In the same graph, draw 3 lines representing ei, e2 and e
Answer #1

Solution-

After running the regression analysis in the excel, we obtain the following regression output-

SUMMARY OUTPUT Regression Statistics Multiple R R Square 0.951612903 Adjusted R Square0.811134235 Standard Error 4.839354796

A. Slope = 1.90323

B. It means that when value of X increases by 1, then value of Y increases by 1.90323

C. Regression equation is-

Y = -1.419 + 1.90323 * X

D. Yes it would be appropriate when X=10 to be calculated using model.

When X = 10, then Y = -1.419 + 1.90323 * 10

= 17.6133

E. e1 = Actual value - predicted value

= -2 - ( -1.419 + 1.90323 * 1)

= -2.484

Answers

TY!